3 -2 3 Find a nonsingular matrix P such that P-1AP is diagonal where A = 0 3 -2 0-3 2
Answer #1

A= 3-2 3 0 - 3 2 NOW, detCA-x1) = 0 gires. 3-2 - 2 3. 3 -2 2ral (3-2) {6–5x+x²-6} → (3-X). X (X-5) = 0 x = 0, 3, 5 are 0,3,5Let O D arhich is a diagonal matrix o 5 5 5 and invertible matrix. is an 6 o 2 5 NOW, -2 AP= 6 3 - 2 3 3 -2 3 2 9 3 25 O O 10